Barton Industries expects that its target capital structure for raising funds in the future for its capital budget will consist of 40% debt, 5% preferred stock, and 55% common equity. Note that the firm's marginal tax rate is 25%. Assume that the firm's cost of debt, rd, is 7.7%, the firm's cost of preferred stock, rps, is 7.2% and the firm's cost of equity is 11.7% for old equity, rs, and 12.3% for new equity, re. What is the firm's weighted average cost of capital (WACC1) if it uses retained earnings as its source of common equity? Do not round intermediate calculations. Round your answer to three decimal places.

  %

What is the firm’s weighted average cost of capital (WACC2) if it has to issue new common stock? Do not round intermediate calculations. Round your answer to three decimal places.

  %
Answer #1

After-tax cost of debt=7.7*(1-tax rate)

=7.7*(1-0.25)=5.775%

WACC=Respective costs*Respective weight

a.WACC=(5.775*0.4)+(7.2*0.05)+(11.7*0.55)

=9.105%

b.WACC=(5.775*0.4)+(7.2*0.05)+(12.3*0.55)

=9.435%